# Benchmarking Fake Voice Detection in the Fake Voice Generation Arms Race

### v1

@@ 3 Fake Voice Detectors
To advance the field of fake voice detection, numerous studies [68, 84] have investigated fake voice detection techniques, which can be broadly classified into two categories: pipeline and end-to-end detectors. While pipeline approaches separate feature extraction and classification, End-to-End (E2E) models operate directly on raw audio waveforms. The following section introduces both frameworks in detail.
@@ 3.1 Feature Extraction
As the front-end and a crucial component of pipeline-based detection systems, feature extraction aims to derive discriminative representations by capturing artifacts indicative of synthetic or manipulated speech. Depending on the feature generation method, existing approaches can be broadly divided into three categories: hand-crafted spectral features, deep learning (DL) features, and other analysis techniques.
@@ 3.1.1 Hand-Crafted Spectral Features
Hand-crafted features are based on human-interpretable properties of audio and leverage domain expertise to focus on characteristics known to be effective in distinguishing between real and fake voices. These features have proven to be a strong baseline in fake voice detection, offering a reliable foundation for capturing discriminative artifact patterns. Among these, spectral features describe the properties of an audio signal in the frequency domain, typically computed from the power spectrum or magnitude spectrum of the signal, which illustrate how a signal’s energy is distributed across frequencies. Notable spectral features, including the Mel Frequency Cepstral Coefficient (MFCC) [22] and the Linear Frequency Cepstral Coefficient (LFCC) [3], are particularly effective in capturing distinct patterns indicative of fake voices.
A spectrogram is a two-dimensional visual representation that illustrates how the frequency content of a signal evolves over time, typically generated by applying the Short-Time Fourier Transform (STFT) to overlapping segments of the audio signal. Among various spectrogram types, the Mel-Spectrogram and Constant-Q Transform (CQT) Spectrogram [28, 20, 92] are the most commonly used in speech processing due to their effectiveness in capturing relevant acoustic features. After extracting spectral features, supervised learning models such as Residual Networks (ResNet) and Recurrent Neural Networks (RNNs) are often employed to generate deep embeddings [51, 30]. For example, [11] utilize an U-Net structure to extract features from CQT-Spectrogram; while in [16] LFCC features and ECAPA-TDNN are combined to generate embeddings.
@@ 3.1.2 Deep Learning Features
Due to recent progress in deep learning (DL), neural network architectures are now employed to derive learnable embeddings that represent the essential traits of raw speech, complementing conventional hand-crafted features. While early approaches relied on traditional machine learning techniques and convolutional neural networks (CNNs), the field has evolved to incorporate more advanced architectures such as graph neural networks (GNNs), Transformers, and other modern deep learning models.
In recent years, self-supervised learning (SSL) frameworks have gained significant traction for feature extraction in fake voice detection, exhibiting remarkable advantages over conventional acoustic and supervised learning-based representations [75, 114]. Xie et al. [138] demonstrate the notable efficacy of Wav2Vec features when implemented within a Siamese network framework. Based on this study, Tak et al. [114] further explore the application of Wav2Vec-2.0 as a front-end feature extractor, combined with a spectro-temporal graph attention network (AASIST) as the back-end classification. Similarly, Guo et al. [37] integrate the WavLM model with a Multi-Fusion Attentive classification to enhance fake speech detection capabilities. In contrast, Wang et al. [126] employ embeddings from the pre-trained HuBERT as inputs.
@@ 3.2 Classification
Traditional back-end classifiers for fake speech detection include conventional models such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) [98, 124]. With the advancements in deep learning, convolutional neural networks (CNNs) have gained prominence for their ability to extract localized and hierarchical patterns within speech signals. Wu et al. [137] and Lavrentyeva et al. [63] implemented the Light CNN (LCNN) architecture [136] for synthetic voice identification, incorporating convolutional and max-pooling operations with Max-Feature-Map (MFM) activation to enhance discriminative feature selection.
@@
Zhang et al. [148] introduce XLSR (Wav2Vec-2.0) with SLS Classifier, which employs a Sensitive Layer Selection module to extract discriminative features from transformer layers for fake voice detection. XLSR-Conformer [100] combines XLS-R with conformer architecture using classification tokens for variable-length utterances. Further refinement through XLSR-Conformer with TCM [120] incorporates Temporal-Channel Modeling to capture dependencies in audio signals, enhancing fake voice detection.
@@ 3.3 E2E Model
End-to-end approaches, which operate directly on raw audio waveforms without requiring explicit feature extraction, have gained significant attention in voice-related tasks, such as fake voice detection. Among these, RawNet2 [46] is a neural network originally designed for speaker verification and identification. It processes raw speech signals using a sinc convolution layer for initial feature extraction, followed by multiple residual blocks to learn deeper representations. Tak et al. [113] were the first to adapt RawNet2 for anti-spoofing tasks. Building on this, Wang et al. [131] proposed a joint optimization strategy leveraging a weighted additive angular margin loss, further improving the performance of the RawNet2-based detection system.
Another notable architecture, SincNet [91], is also designed to process raw audio directly, leveraging parameterized sinc functions to efficiently learn relevant frequency bands. Ge et al. [31] employed sinc-shaped filters and partial channel connections for efficiency. Zeinali et al. [146] introduce a fusion framework that combines SincNet and VGG architectures to enhance fake voice detection performance. Additionally, Huang et al. [40] propose DFSincNet, an end-to-end model that enhances SincNet by incorporating discriminative frequency information, with a particular focus on high-frequency components to improve spoofing detection accuracy. RawBMamba [18] utilizes sinc and convolutional layers for short-range features and bidirectional mamba for long-range features, with a fusion module integrating embeddings to enhance audio representation.
@@ 4 Cross-domain Evaluation
To address the limitations of common fake voice detector evaluations, which typically rely on static and aggregated datasets, thereby obscuring the unique artifacts introduced by different fake voice generators, we propose a cross-domain, comprehensive assessment. This evaluation includes two experiments involving 22 types of fake voices, generated using 15 distinct voice synthesis models from three major categories: traditional pipeline models, neural codec language models, and end-to-end models, along with five existing datasets for comparison. To assess the robustness and generalizability of detection systems, we evaluate these voices using 8 different fake voice detectors, analyzing their performance on both individually generated voices and aggregated datasets. Since detector performance can vary significantly depending on the type of fake voice, with some detectors excelling in specific scenarios while failing in others, our approach provides a more nuanced understanding of their vulnerabilities. Additionally, we introduce two metric systems designed to quantify the overall effectiveness of both fake voice generators and detectors from multiple perspectives, enabling fair comparison and ranking across different systems.
@@ 4.1 Experimental Design
Fake Voices Selection 15 state-of-the-art fake voice generators were selected covering all categories introduced in Section 2, including traditional pipeline models, codec language models, and end-to-end architectures. As traditional TTS systems involve vocoders for converting Mel-spectrograms to waveforms, we also evaluated standalone vocoder performance for audio reconstruction. To ensure comprehensive assessment, we included public datasets containing both real and fake audio: the latest ASVspoof-2024, ASVspoof-2021 LA and DF, the Fake or Real (FoR) dataset, and the CFAD dataset for bilingual analysis. A detailed overview of selected fake voices is presented in Table 2.
@@
Fake Voice Generation Process LibriSpeech is a widely used open-source dataset designed for automatic speech recognition (ASR) and text-to-speech (TTS) research. It contains English speech derived from audiobooks, paired with aligned text transcriptions. In our experiments, we used the test-clean subset, which includes 8.56 hours of audio from 39 speakers, sampled at 24 kHz in a multi-speaker English corpus. For TTS systems, we used the transcriptions as input to generate fake (synthetic) voices. The original LibriSpeech audio served as the reference for evaluating the similarity and authenticity of the generated speech. For TTS systems with zero-shot capabilities, able to mimic specific speaker characteristics, the generated fake voices closely resembled the original speaker. In contrast, TTS models without zero-shot functionality produced fake voices using a default or generic speaker profile. For vocoder evaluation (audio reconstruction), we first converted the original waveforms into Mel-spectrograms and then back into waveforms using vocoders, treating the output as synthetic speech. We compared these reconstructed outputs with the original waveforms to evaluate the vocoder’s performance. All TTS models, vocoders, and inference pipelines used in our experiments were sourced from publicly available pre-trained models on Hugging Face.
Detector Selection & Implementation Similar to the approach for selecting fake voice generators, we carefully chose 8 state-of-the-art fake voice detectors for evaluation. Our selection criteria focused on accessibility and benchmark performance on the ASVspoof dataset, which is widely used in the field for evaluating fake voice detection systems. We selected two main types of detectors: end-to-end (E2E) models and advanced models that use transformer-based front-end for feature extraction. While the latter type of detector may share a similar front-end, their back-end classifiers differ significantly, resulting in varied performance, particularly in cross-domain evaluations. The selected detectors are listed in Table 3which we divide into two categories: end-to-end detectors and detectors that used XLSR (Wav2Vec-2.0) as feature extractions. We used the official implementations of all eight detectors, each trained on the ASVspoof19-LA dataset.
@@ 4.2.1 Performance Metric for Generator
The Mean Opinion Score (MOS) has been the standard for decades; yet, recent text-to-speech (TTS) systems often exceed the discriminative resolution of its five-point scale and suffer from listener bias, high variance, and costly test administration. [65] Consequently, we adopted a set of objective metrics for TTS and vocoder evaluation that (i) correlate with perceptual quality, (ii) cover intelligibility, naturalness, and similarity, and (iii) computational overhead. Then, we computed an overall score based on those criteria. Exact metric computation procedures are given in Appendix G.
For TTS systems (including Zero-Shot) :
PSNR [2]: Compares maximum signal power to noise power.
DNSMOS[93]: Using deep learning methods to predict subjective audio perceptual quality.
Real-time Factor (RTF): The ratio of the time taken by the model to process or generate one second of audio to the actual duration of that audio; values less than 1 indicate faster-than-real-time performance.
Parameters: The total number of trainable parameters in the model, reflecting its size, complexity, and memory requirements, measured in million.
WER & CER: Evaluate word error rate (WER) and character error rates (CER) using Whisper[90], demonstrating intelligibility.
SIM: WavLM-based [14] speaker similarity between original and generated speech pairs.
For Vocoder models:
PESQ [99]: Perceptual differences between generated and reference speech.
SSIM[132]: Frequency-domain similarity between generated and real spectrograms.
LS-MSE[2]: Mean square error between ground truth and generated log-mel spectrograms.
@@ 4.2.2 Individual Metrics Analysis and Generator Overall Performance
Generator Overall Score Formulation. We define the overall performance of a generator as a weighted aggregation of normalized metrics. For a given task, the composite score is expressed as:
@@
where $M_{i,\text{norm}}$ and $M_{j,\text{norm}}$ denote the $i$-th and $j$-th evaluation metrics, each normalized to the range $(0,1]$. The coefficients $\alpha_{i}$ and $\beta_{j}$ are non-negative weights that satisfy $\sum_{i}\alpha_{i}=1$ and $\sum_{j}\beta_{j}=1$, respectively.
The specific choice of metrics (DNSMOS, SIM, PESQ, PSNR, SSIM, WER, RTF, and model size) and their associated weights are detailed in Appendix G.
Table 4: Performance evaluation of various TTS generators. Column best values highlighted blue, worst red. WER and CER are shown in %. Params = Parameters in Million, Overall = Overall score.
@@
Table 5: Performance evaluation of various Vocoders (audio reconstruction). Column best values highlighted blue, worst red. Params = Parameters in Million, Overall = Overall score.
@@
Overall Performance Results. Flow Matching approaches (MatchaTTS + HiFi-GAN) perform commendably with good DNSMOS scores. For ZSVC systems, YourTTS stands out with high overall score and competitive PSNR/SIM performance while maintaining moderate parameter count (87M). Neural Codec models like FireRedTTS-1S and MaskGCT-TTS deliver strong results but with substantially larger parameter counts (1294M and 2226M respectively). For vocoders, architecture-improved GANs such as BigVGAN and Vocos demonstrate leading performance, while diffusion-based models generally exhibit lower overall scores despite having fewer parameters.

### v2

@@ 3 Fake Voice Detectors
To advance the field of fake voice detection, numerous studies [25, 24] have investigated fake voice detection techniques, which can be broadly classified into two categories: pipeline and end-to-end detectors. While pipeline approaches separate feature extraction and classification, End-to-End (E2E) models operate directly on raw voice waveforms. The following section introduces both frameworks in detail.
@@ 3.1 Feature Extraction
As the front-end and a crucial component of pipeline-based detection systems, feature extraction aims to derive discriminative representations by capturing artifacts indicative of fake voice. Depending on the feature generation method, existing approaches can be broadly divided into three categories: hand-crafted spectral features, Deep Learning (DL) features, and other analysis techniques.
@@ 3.1.1 Hand-Crafted Spectral Features
Hand-crafted spectral features leverage domain expertise to analyze a voice signal’s properties in the frequency domain, providing a strong baseline for fake voice detection. These features, computed from the signal’s power or magnitude spectrum, capture how its energy is distributed across different frequencies.
Prominent examples include Mel Frequency Cepstral Coefficients (MFCC) [62], Linear Frequency Cepstral Coefficients (LFCC) [63], and two-dimensional spectrograms like the Mel-Spectrogram and Constant-Q Transform (CQT) Spectrogram [64, 65]. These representations are effective at highlighting discriminative artifacts indicative of fake speech. Subsequently, deep learning models such as Residual Networks (ResNet), Recurrent Neural Networks (RNNs), or U-Nets are often employed to process these features and generate deep embeddings [66, 67].
A primary challenge, however, is robustness to real-world conditions. Empirical evidence from ASVspoof evaluations shows that performance can be significantly degraded by channel and device variability in physical access attacks [68, 69] and by codec artifacts in logical-access scenarios involving VoIP transmission [70].
@@ 3.1.2 Deep Learning Features
Due to progress in Deep Learning (DL), neural network architectures are now employed to derive learnable embeddings that represent the essential traits of raw voice, complementing conventional hand-crafted features. While early approaches relied on traditional machine learning techniques and convolutional neural networks (CNNs), the field has evolved to incorporate more advanced architectures such as graph neural networks (GNNs), Transformers, and other modern deep learning models.
In recent years, self-supervised learning (SSL) frameworks have gained significant traction for feature extraction in fake voice detection, exhibiting remarkable advantages over conventional acoustic and supervised learning-based representations [71, 72]. Xie et al. [73] demonstrate the notable efficacy of Wav2Vec features when implemented within a Siamese network framework. Based on this study, Tak et al. [72] further explore the application of Wav2Vec-2.0 as a front-end feature extractor, combined with a spectro-temporal graph attention network (AASIST) as the back-end classification. Similarly, Guo et al. [74] integrate the WavLM model with a Multi-Fusion Attentive classification to enhance fake voice detection capabilities. In contrast, Wang et al. [75] employ embeddings from the pre-trained HuBERT as inputs. Despite their accuracy, SSL embeddings and DL front-ends are vulnerable to adversarial perturbations and transfer-based black-box attacks that can evade state-of-the-art audio deepfake detectors, underscoring the need for attack-aware training and evaluation [76, 77, 78]. This risk is especially salient in fraud/social-engineering use cases, where attackers can query or approximate deployed models [79].
@@ 3.2 Classification
Traditional back-end classifiers for fake voice detection include conventional models such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) [80, 81]. With the advancements in deep learning, convolutional neural networks (CNNs) have gained prominence for their ability to extract localized and hierarchical patterns within voice signals. Wu et al. [82] and Lavrentyeva et al. [83] implemented the Light CNN (LCNN) architecture [84] for synthetic voice identification, incorporating convolutional and max-pooling operations with Max-Feature-Map (MFM) activation to enhance discriminative feature selection.
@@
Zhang et al. [92] introduce XLSR (Wav2Vec-2.0) with SLS Classifier, which employs a Sensitive Layer Selection module to extract discriminative features from transformer layers for fake voice detection. XLSR-Conformer [93] combines XLS-R with conformer architecture using classification tokens for variable-length utterances. Further refinement through XLSR-Conformer with TCM [94] incorporates Temporal-Channel Modeling to capture dependencies in voice signals, enhancing fake voice detection. In deployments such as telephony or VoIP, smart speakers, back-ends must generalize to unseen spoofing methods and environmental shifts; ASVspoof analyses consistently highlight generalization gaps from lab conditions to field conditions, motivating cross-corpus and cross-attack evaluation [95, 96]. Classifiers should therefore be assessed not only on raw synthetic speech but also under compression, re-recording, and laundering processes that attackers use to obfuscate artifacts [97].
@@ 3.3 E2E Model
End-to-end approaches, which operate directly on raw voice waveforms without requiring explicit feature extraction, have gained significant attention in voice-related tasks, such as fake voice detection. Among these, RawNet2 [98] is a neural network originally designed for speaker verification and identification. It processes raw voice signals using a sinc convolution layer for initial feature extraction, followed by multiple residual blocks to learn deeper representations. Tak et al. [99] were the first to adapt RawNet2 for anti-spoofing tasks. Building on this, Wang et al. [100] proposed a joint optimization strategy leveraging a weighted additive angular margin loss, further improving the performance of the RawNet2-based detection.
Another notable architecture, SincNet [101], is also designed to process raw voice directly, leveraging parameterized sinc functions to efficiently learn relevant frequency bands. Ge et al. [1] employed sinc-shaped filters and partial channel connections for efficiency. Zeinali et al. [102] introduce a fusion framework that combines SincNet and VGG architectures to enhance fake voice detection performance. Additionally, Huang et al. [103] propose DFSincNet, an end-to-end model that enhances SincNet by incorporating discriminative frequency information, with a particular focus on high-frequency components to improve spoofing detection accuracy. RawBMamba [104] utilizes sinc and convolutional layers for short-range features and bidirectional mamba for long-range features, with a fusion module integrating embeddings to enhance voice representation. E2E detectors must be evaluated against adaptive adversaries and post-processing laundering (e.g., resampling, compression, playback–re-recording): recent benchmarks show notable performance drops when synthetic audio is intentionally processed to evade detection, and gradient-based attacks can reduce margins [97, 105].
@@ 4 Cross-domain Evaluation
To address the limitations of common fake voice detector evaluations, which typically rely on static and aggregated datasets, thereby obscuring the unique artifacts introduced by different fake voice generators, we propose a cross-domain, comprehensive assessment. This evaluation includes two experiments involving 17 types of fake voices, generated using distinct fake voice generator from three major categories: traditional pipeline models, neural codec language models, and end-to-end models, along with five existing datasets for comparison. To assess the robustness and generalizability of detection systems, we evaluate these voices using 8 different fake voice detectors, analyzing their performance on both individually generated voices and aggregated datasets. Since detector performance can vary significantly depending on the type of fake voice, with some detectors excelling in specific scenarios while failing in others, our approach provides a more nuanced understanding of their vulnerabilities. Additionally, we introduce two metric systems designed to quantify the overall effectiveness of both fake voice generators and detectors from multiple perspectives, enabling fair comparison and ranking across different systems.
@@ 4.1 Experimental Design
4.1.1 Threat Model
  • Adversary’s Goal: To create fake voice that can successfully bypass automated fake voice detection systems, such as those used as countermeasures for ASV. The attacks include both zero-shot voice cloning and the generation of non-cloned but highly natural fake voice.
  • Adversary’s Capabilities: The adversary is assumed to have access to publicly available, State-Of-The-Art (SOTA) open-source fake voice generation models. They do not have white-box access to the target detectors but aim to create voice that is broadly evasive. This models a realistic scenario where sophisticated tools are readily accessible to a wide range of actors.
4.1.2 Selection Criteria
Fake Voice Generators and Datasets: Our selection of 15 generators was guided by three criteria: (1) Architectural Diversity, ensuring coverage of traditional pipelines, end-to-end models, and modern neural codec architectures as detailed in Section 2; (2) SOTA Performance, including models recognized for producing high-fidelity output; and (3) Public Accessibility, using open-source models with pre-trained weights to ensure reproducibility. For existing datasets, we selected widely-used benchmarks (ASVspoof series, FoR) to ground our results in established evaluation paradigms. We included the CFAD dataset to test cross-lingual generalization. While other datasets like ADD [28] exist, our evaluation focuses primarily on the English language. A detailed overview is in Table 1.
@@ 4.1.2 Selection Criteria
Fake Voice Detectors: The 8 detectors were chosen based on: (1) Strong Performance on the public ASVspoof benchmark; (2) Architectural Variety, including both end-to-end models and those using powerful pre-trained front-ends like Wav2Vec-2.0; and (3) Availability of official open-source implementations. This ensures our evaluation reflects the current capabilities of leading detection paradigms. All selected detectors (Table 2) were used with their official implementations, trained on the ASVspoof19-LA dataset.
@@
Fake Voice Generation Process. LibriSpeech is a widely used open-source dataset designed for automatic speech recognition (ASR) and text-to-speech (TTS) research. It contains English voice derived from audiobooks, paired with aligned text transcriptions. In our experiments, we used the test-clean subset, which includes 8.56 hours of voice from 39 speakers, sampled at 24 kHz in a multi-speaker English corpus. For TTS systems, we used the transcriptions as input to generate fake voices. The original LibriSpeech voice served as the naturalistic reference. For TTS systems with zero-shot capabilities, able to mimic specific speaker characteristics, the generated fake voices closely resembled the original speaker. In contrast, TTS models without zero-shot functionality produced fake voices using a default or generic speaker profile. For audio reconstruction, we first converted the original waveforms into Mel-spectrograms and then back into waveforms using vocoders, treating the output as fake voice. We compared these reconstructed outputs with the original waveforms to evaluate the performance. All inference pipelines used in our experiments were sourced from publicly available pre-trained models on Hugging Face.
@@ 4.2.1 Performance Metric for Generator
The Mean Opinion Score (MOS) has been the standard for decades; yet, recent Text-to-Speech (TTS) systems often exceed the discriminative resolution of its five-point scale and suffer from listener bias, high variance, and costly test administration. [107] Consequently, we adopted a set of objective metrics for TTS and vocoder evaluation that (i) correlate with perceptual quality, (ii) cover intelligibility, naturalness, and similarity, and (iii) computational overhead. Then, we computed an overall score based on those criteria.
The performance of Text-to-Speech (TTS) systems, including zero-shot models, is evaluated using several metrics. PSNR [22] compares the maximum signal power to noise power, while DNSMOS [108] uses deep learning to predict subjective audio perceptual quality. The Real-time Factor (RTF) measures efficiency, defined as the ratio of processing time to the audio’s duration, where values under 1 signify faster-than-real-time performance. Model size and complexity are indicated by the total number of trainable Parameters. Intelligibility is assessed through WER (word error rate) and CER (character error rate) using Whisper [109], and speaker similarity (SIM) is measured with a WavLM-based model [110] between original and generated voice pairs. For vocoder models specifically, evaluation includes PESQ [111] to quantify perceptual differences, SSIM [112] to assess frequency-domain similarity between spectrograms, and LS-MSE [22] to calculate the mean square error between the ground truth and generated log-mel spectrograms.
@@ 4.2.2 Individual Metrics Analysis and Generator Overall Performance
Generator Overall Score Formulation. We compute an overall score as a weighted sum of normalized metrics:
@@
where $M_{\cdot,\text{norm}}\in(0,1]$ are per-metric values normalized as detailed in Appendix.
For TTS, we assign highest mass to perceptual naturalness and speaker identity fidelity: DNSMOS is a no-reference proxy shown to correlate with human MOS for speech quality, while SIM captures speaker similarity, a standard proxy for identity preservation in zero-shot TTS with voice conversion. Intelligibility via WER is a long-standing ASR metric that tracks content accuracy. Fidelity metrics (PSNR) and structure-aware SSIM remain informative for reconstruction artifacts. Finally, deployment efficiency matters but should not dominate perceptual outcomes, so we give small weights to latency (RTF) and model footprint (LogParams).
For Audio Reconstruction, where speaker identity is typically not a target constraint, we shift mass to full-reference quality: DNSMOS complements reference-based measures; PESQ is the ITU-T standard for perceived speech quality; SSIM captures structural fidelity beyond pixelwise error; PSNR remains a widely used baseline; and efficiency terms, including RTF and LogParams, ensure practical deployability. This scheme prioritizes what end users perceive (naturalness or identity for TTS and perceptual fidelity for reconstruction) and includes efficiency as a secondary objective.
Table 3: Performance evaluation of various TTS generators. The best and worst performing values for each metric are highlighted in blue and red, respectively. WER and CER are shown in %. Params = Parameters in Million, Overall = Overall score.
@@
Table 4: Performance evaluation of various Vocoders (audio reconstruction). The best and worst performing values for each metric are highlighted in blue and red, respectively. Params = Parameters in Million, Overall = Overall score.
@@
Overall Performance Results. Flow Matching approaches (MatchaTTS + HiFi-GAN) perform commendably with good DNSMOS scores. For ZSVC systems, YourTTS stands out with high overall score and competitive PSNR/SIM performance while maintaining moderate parameter count (87M). Neural Codec models like FireRedTTS-1S and MaskGCT-TTS deliver strong results but with substantially larger parameter counts (1294M and 2226M respectively). For vocoders, architecture-improved GANs such as BigVGAN and Vocos demonstrate leading performance, while diffusion-based generally exhibit lower overall scores despite fewer parameters.